OpenStudy (anonymous):

Which of the following is false for an exponential growth function y = C(1 + r)t? when C > 0, the function decreases from left to right the function has a y-intercept of (0, C) the graph of this function will be identical to the graph of y = abx, if a = c and b = 1 + r if C > 0 and r > 0, then the graph will only have positive y-values.

OpenStudy (phi):

you should use ^ to show exponents like this y = C (1+r)^t

OpenStudy (anonymous):

That doesn't really help with my question, but yeah I know, thanks.

OpenStudy (phi):

you could try a few "test cases" to test when C > 0, the function decreases from left to right pick C=1 , and r=1 y = C(1 + r)^t becomes y = 2^t t=0 gives 2^0=1 t=1 gives 2^1 =2 so y goes from 1 to 2 as t goes from 0 to 1 when C > 0, the function decreases from left to right is the function decreasing ?

OpenStudy (phi):

That doesn't really help with my question, It helps people decipher the question.